# Suicide Mortality in Spain (2010-2022): Temporal Trends, Spatial Patterns, and Risk Factors

### v1

@@ 4.1 Analysis of sex-specific temporal and spatial patterns
Table 3 presents the Deviance Information Criterion (DIC) and Watanabe-Akaike Information Criterion (WAIC) values used to assess and compare the fit of the age-time and age-space models separately for males and females. In line with the descriptive analysis, models including interaction effects consistently provide a better fit than their additive counterparts in all cases. For age-time models, the Type II interaction model provides the best fit for male suicide rates according to the model selection criteria. This suggests that, beyond the main structured effects, the age-specific evolution during the study period shows a temporal pattern that does not necessarily have to be similar across consecutive age groups. In contrast, for female suicide rates we select the model incorporating a fully structured interaction random effect (i.e., Type IV model). For the age-space models, we select the Type III interaction for both male and female data, indicating that the interaction effect captures non-spatial variability across provinces while preserving correlation between adjacent age groups.
Table 3: Comparison of age-time and age-space interaction models by sex, based on DIC and WAIC criteria.
@@
Figure 4: Posterior median estimates and 95% credible intervals of sex-specific age (left figure) and time (right figure) patterns of suicide mortality rates.
@@
Figure 5: Posterior median estimates (left figures) and exceedence probabilities (right figures) of sex-specific spatial patterns of suicide mortality rates.
@@
The temporal evolution of sex-specific suicide mortality rate estimates by age group is presented in Figure 6. As expected, the credible intervals for female estimates are substantially wider, reflecting that the number of recorded deaths is approximately three times higher in males than in females across most age groups. In general, suicide mortality rates progressively increase with age, except among males aged 40-70 and females over 50, where no significant differences are observed. Over the study period, temporal trends shows a slight decline among elderly males, whereas for females in these age groups (and in most others) the trend clearly increases.
@@
Figure 6: Temporal evolution of posterior median estimates and 95% credible intervals of suicide mortality rates by age group. Different scales are used for the Y axis to improve the visualization of the rates.
Finally, the maps illustrating the geographical distribution of suicide mortality rates by age group are shown in Figure 7 (males) and Figure 8 (females). In both cases, the maps reveal a consistent geographical pattern across age groups, with overlapping areas of high and low rates. Nonetheless, some differences emerge between specific age groups. Among the male population, provinces in western and southern inland Spain (including Galicia, much of Castilla-La Mancha, and the interior of Andalusia) consistently show higher mortality rates in the older age groups, particularly from 50-59 years onward, with rates especially pronounced in the 70-79 and 80+ age groups. In contrast, some provinces in the north, northeast, and along the Mediterranean coast (such as the Basque Country, Navarra, Aragón, and the Valencian Community) tend to exhibit lower mortality rates within the same older age groups. Among the female population, the underlying geographic pattern exhibits stronger spatial dependence than in males. In particular, most regions in the northwest of Spain show consistently high mortality rates across all age groups. Meanwhile, some northern provinces (such as Navarra and Aragón) also display relatively high mortality rates in younger age groups, but a downward trend emerges as age increases. In contrast, certain southern regions exhibit the opposite pattern: mortality rates there not only remain above the national average but also show a clear increasing trend among older age groups.
@@
Figure 7: Maps of posterior median estimates of male suicide mortality rates by age group.
@@
Figure 8: Maps of posterior median estimates of female suicide mortality rates by age group.
@@ 4.2 Analysis of socioeconomic and geographic disparities
Figure 9: Spatial distribution of percentage of the population living in rural areas (left) and unemployment rate for males (center) and females (right).
Figure 10: Sex specific temporal trends in the annual unemployment rate (left) and the at-risk-of-poverty rate (right).
@@ 5 Discussion
This study provides a comprehensive analysis of suicide mortality in Spain, highlighting notable differences across sex, age group, and geographic region. Although male suicide rates consistently exceeded those of females between 2010 and 2022, we observed a pronounced upward trend among females. This increase was particularly evident in younger age groups, with posterior median estimates indicating a 173% rise among those aged 10-19 and a 124% rise among those aged 20-29, as well as a 54% increase among females aged 50-59. In terms of spatial patterns, males exhibited greater heterogeneity across provinces, with a larger number of areas showing significantly elevated rates. In contrast, females displayed a smoother spatial distribution. Nonetheless, two regions -Galicia in the northwest and southern Andalusia- were consistently identified in both sexes as having mortality rates markedly above the national mean. Additionally, Navarra and its neighboring eastern provinces displayed particularly elevated suicide rates among younger females.
@@
When examining the influence of socioeconomic and contextual factors, we found that suicide rates were positively associated with the proportion of the population living in rural areas among males, and with the annual unemployment rate among females. In contrast, provincial unemployment and at-risk-of-poverty rates showed no clear association with suicide mortality in either sex. The observed association between male suicide rates and rurality is consistent with previous studies conducted in Ireland (O’Farrell et al., , 2016), Germany (Helbich et al., , 2017), the USA (Fontanella et al., , 2015; Ivey-Stephenson, , 2017) and England (Middleton et al., , 2006; Gascoigne et al., , 2025), among others. This pattern is further supported by a recent systematic review and meta-analysis (Davico et al., , 2025), which demonstrated that lower population density, commonly used as a proxy for rurality, was associated with higher suicide rates in the majority of countries examined.
@@
Considering the multifactorial nature of suicide risk, arising from the interplay of biological, psychological, clinical, and socio-environmental factors (Turecki et al., , 2019), Bayesian hierarchical models with structured random effects offer a powerful framework to capture complex spatial, temporal, and age-related patterns in mortality rates. These models enable comprehensive uncertainty quantification and incorporate correlations across provinces, time periods, and age groups, resulting in more reliable and robust estimates. Another key strength of our work is the use of the simplified spatial+ approach within an ecological regression framework, which accounts for potential spatial and temporal confounding between fixed and random effects, thereby reducing the risk of biased fixed-effect estimates that could result if these dependencies were ignored.
@@
One limitation of our study is that attempts to fit fully stratified age-space-time models led to excessive smoothing. This is because the data was too sparse at this level of resolution, which caused excessive smoothing and could have led to incorrect interpretations (Retegui et al., , 2025). Furthermore, the absence of monthly suicide mortality data at the subnational level prevented an examination of seasonal patterns, such as the peak in suicide rates during spring and summer that has been previously documented in Spain (Salmeron et al., , 2013). Finally, although some recent studies have attempted to explore the impact of the COVID-19 pandemic on suicide mortality (Fouillet et al., , 2023; Ramalle-Gómara et al., , 2023), the span of our study from 2010 to 2022 makes it difficult to isolate and accurately assess these specific effects.
Acknowledgments
@@
This work has been supported by projects PID2024-155382OB-I00 (funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE), PID2020-113125RB-I00/MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation), and UNEDPAM/PI/PR24/01A (Centro Asociado UNED - Pamplona).

### v2

@@ 4.2 Analysis of sex-specific temporal and spatial patterns
Table 3 presents the Deviance Information Criterion (DIC) and Watanabe-Akaike Information Criterion (WAIC) values used to assess and compare the fit of the age-time and age-space models separately for males and females. Both DIC and WAIC are measures of model fit that balance goodness-of-fit and model complexity: lower values indicate a better trade-off between how well the model explains the data and the effective number of parameters. In Table 3, the smallest values are highlighted in bold to indicate the preferred model for each scenario. In line with the descriptive analysis, models including interaction effects consistently provide a better fit than their additive counterparts in all cases. For age-time models, the Type II interaction model provides the best fit for male suicide rates according to the model selection criteria. This suggests that, beyond the main structured effects, the age-specific evolution during the study period shows a temporal pattern that does not necessarily have to be similar across consecutive age groups. In contrast, for female suicide rates we select the model incorporating a fully structured interaction random effect (i.e., Type IV model). For the age-space models, we select the Type III interaction for both male and female data, indicating that the interaction effect captures non-spatial variability across provinces while preserving correlation between adjacent age groups.
Table 3: Comparison of age-time and age-space interaction models by sex, based on DIC and WAIC criteria. The smallest values are highlighted in bold to indicate the preferred model for each scenario.
@@
Figure 4: Posterior median estimates and 95% credible intervals of sex-specific age (a) and time (b) patterns of suicide mortality rates.
@@
Figure 5: Sex-specific spatial patterns of suicide mortality rates. (a) Posterior median estimates for males; (b) Posterior exceedence probabilities for males; (c) Posterior median estimates for females; (d) Posterior exceedence probabilities for females.
@@
The temporal evolution of sex-specific suicide mortality rate estimates by age group is presented in Figure 6. As expected, the credible intervals for female estimates are substantially wider, reflecting that the number of recorded deaths is approximately three times higher in males than in females across most age groups. In general, suicide mortality rates progressively increase with age, except among males aged 40–70 and females over 50, where no significant differences are observed. Over the study period, temporal trends shows a slight decline among elderly males, whereas for females in these age groups (and in most others) the trend clearly increases.
@@
Figure 6: Temporal evolution of posterior median estimates and 95% credible intervals of suicide mortality rates by age group. Different scales are used for the Y axis to improve the visualization of the rates. (a) Male population; (b) Female population.
@@
Figure 7: Maps of posterior median estimates of male suicide mortality rates by age group. The breaks in the color scales were defined using the quintiles of the estimated rates within each age group. (a) Age group 10–19; (b) Age group 20–29; (c) Age group 30–39; (d) Age group 40–49; (e) Age group 50–59; (f) Age group 60–69; (g) Age group 70–79; (h) Age group 80+.
@@
Figure 8: Maps of posterior median estimates of female suicide mortality rates by age group. The breaks in the color scales were defined using the quintiles of the estimated rates within each age group. (a) Age group 10–19; (b) Age group 20–29; (c) Age group 30–39; (d) Age group 40–49; (e) Age group 50–59; (f) Age group 60–69; (g) Age group 70–79; (h) Age group 80+.
Finally, the maps illustrating the geographical distribution of suicide mortality rates by age group are shown in Figure 7 (males) and Figure 8 (females). For visualization purposes, the breaks in the color scale were defined using the quintiles of the estimated rates within each age group, so comparisons across panels should be interpreted with caution. In both cases, the maps reveal a consistent geographical pattern across age groups, with overlapping areas of high and low rates. Nonetheless, some differences emerge between specific age groups. Among the male population, provinces in western and southern inland Spain (including Galicia, much of Castilla-La Mancha, and the interior of Andalusia) consistently show higher mortality rates in the older age groups, particularly from 50–59 years onward, with rates especially pronounced in the 70–79 and 80+ age groups. In contrast, some provinces in the north, northeast, and along the Mediterranean coast (such as the Basque Country, Navarra, Aragón, and the Valencian Community) tend to exhibit lower mortality rates within the same older age groups. Among the female population, the underlying geographic pattern exhibits stronger spatial dependence than in males. In particular, most regions in the northwest of Spain show consistently high mortality rates across all age groups. Meanwhile, some northern provinces (such as Navarra and Aragón) also display relatively high mortality rates in younger age groups, but a downward trend emerges as age increases. In contrast, certain southern regions exhibit the opposite pattern: mortality rates there not only remain above the national average but also show a clear increasing trend among older age groups.
@@ 4.3 Analysis of socioeconomic and geographic disparities
Figure 9: Spatial distribution of percentage of the population living in rural areas (a) and unemployment rate for males (b) and females (c).
Figure 10: Sex specific temporal trends in the annual unemployment rate (a) and the at-risk-of-poverty rate (b).
@@ 5 Discussion
This study provides a comprehensive analysis of suicide mortality in Spain, highlighting notable differences across sex, age group, and geographic region. Our definition of suicide mortality is limited to ICD-10 codes X60–X84. While other codes could potentially be considered suicides (Y87: ’Sequelae or late effects of intentional self-harm, assault, and events of undetermined intent’, Y10–Y19: ’Poisoning of undetermined intent’, and Y20–Y34: ’Other events of undetermined intent’), the exclusion of these codes may result in a very small undercount of suicide deaths. However, their contribution is minimal (accounting for less than 2% of total cases during the study period) and is unlikely to materially affect the spatial or temporal patterns observed in our study.
@@
Although male suicide rates consistently exceeded those of females between 2010 and 2022, we observed a pronounced upward trend among females. This increase was particularly evident in younger age groups, with posterior median estimates indicating a 173% rise among those aged 10–19 and a 124% rise among those aged 20–29, as well as a 54% increase among females aged 50–59. In terms of spatial patterns, males exhibited greater heterogeneity across provinces, with a larger number of areas showing significantly elevated rates. In contrast, females displayed a smoother spatial distribution. Nonetheless, two regions -Galicia in the northwest and southern Andalusia- were consistently identified in both sexes as having mortality rates markedly above the national mean. Additionally, Navarra and its neighboring eastern provinces displayed particularly elevated suicide rates among younger females.
Our findings also show a marked increase in suicide mortality among the oldest male age groups in Spain, particularly from age 80 onwards, a pattern that aligns with evidence reported in the international literature. Previous research has demonstrated that suicide mortality rises progressively with age and reaches its highest levels among older men, including those aged 85-90 years (shah2016suicide). Other studies similarly report especially elevated rates among individuals aged 80 years and older (naghavi2019global), and evidence suggests that this trend may even persist among centenarians (de2022late). Physical illness and functional disability, both highly prevalent in late life, may help explain the particularly elevated suicide rates observed among the oldest men. In Spain, mental health organizations have highlighted unwanted loneliness as an additional risk factor in this population. Physical health problems and functional limitations often entail loss of autonomy, chronic pain, social withdrawal, and depressive symptoms, all of which have been identified as important contributors to suicide risk in older adults (shah2016suicide). Furthermore, some studies suggest that the impact of physical health problems on suicidal behaviour may vary across advanced age groups. For example, a systematic review reported that hospitalization and certain chronic diseases, such as cancer, were significant predictors of suicide among men aged 80 years and older, but not among those aged 70-79 (fassberg2016systematic).
@@
When examining the influence of socioeconomic and contextual factors, we found that suicide rates were positively associated with the proportion of the population living in rural areas among males, and with the annual unemployment rate among females. In contrast, provincial unemployment and at-risk-of-poverty rates showed no clear association with suicide mortality in either sex. The observed association between male suicide rates and rurality is consistent with previous studies conducted in Ireland (o2016area), Germany (helbich2017spatiotemporal), the USA (fontanella2015widening; ivey2017suicide) and England (middleton2006geography; gascoigne2025spatio), among others. This pattern is further supported by a recent systematic review and meta-analysis (davico2025relationship), which demonstrated that lower population density, commonly used as a proxy for rurality, was associated with higher suicide rates in the majority of countries examined. However, the positive association between annual unemployment and female suicide rates is somewhat paradoxical, as previous studies in Spain and other countries have generally reported stronger associations among males (cunningham2022gendered; blazquez2023associations). Interestingly, kposowa2019new observed a similar pattern in the United States, where unemployed women exhibited higher suicide rates than unemployed men, consistent with our findings and suggesting that gender-specific vulnerabilities may underlie this relationship.
@@
Considering the multifactorial nature of suicide risk, arising from the interplay of biological, psychological, clinical, and socio-environmental factors (turecki2019suicide), Bayesian hierarchical models with structured random effects offer a powerful framework to capture complex spatial, temporal, and age-related patterns in mortality rates. These models enable comprehensive uncertainty quantification and incorporate correlations across provinces, time periods, and age groups, resulting in more reliable and robust estimates. Another key strength of our work is the use of the simplified spatial+ approach within an ecological regression framework, which accounts for potential spatial and temporal confounding between fixed and random effects, thereby reducing the risk of biased fixed-effect estimates that could result if these dependencies were ignored.
One limitation of our study is that attempts to fit fully stratified age-space-time models led to excessive smoothing. This is because the data was too sparse at this level of resolution, which caused excessive smoothing and could have led to incorrect interpretations (retegui2025prior). The analysis of suicide mortality is further limited by the potential underreporting and regional inconsistencies in registry data, largely due to the challenging nature of cause-of-death classification. For instance, official suicide rates in Spain have historically been lower than those in some other European countries (snowdon2021spain). Other researchers (SANTURTUN2022150) posit that this discrepancy may be attributed, in part, to the fact that Spain’s Statistical Registry of Deaths with Judicial Intervention was previously filled out by a civil servant, who often lacked adequate information about the cause of death, rather than a forensic doctor.
@@
Furthermore, the absence of monthly suicide mortality data at the subnational level prevented an examination of seasonal patterns, such as the peak in suicide rates during spring and summer that has been previously documented in Spain (salmeron2013time). In addition, publicly available data from the Spanish National Statistics Institute are often not accessible at the level of disaggregation required for our analysis (i.e., by sex, age-group, year, and province). This limited the inclusion of additional potential risk factors, such as socioeconomic level, proportion of migrant population, antidepressant prescription rates, or access to mental health services. Finally, although some recent studies have attempted to explore the impact of the COVID-19 pandemic on suicide mortality or suicide behavior (dube2021suicide; fouillet2023reactive; lantos2024impact; ramalle2023trends; DELATORRELUQUE202429), we find that it is still necessary to enlarge the study period to better understand the real impact of the pandemic. The span of our study from 2010 to 2022 makes it difficult to isolate and accurately assess these specific effects.
6 Conclusion
This study highlights several key conclusions regarding suicide mortality in Spain from 2010 to 2022, emphasizing distinct patterns based on gender and geographical location. A notable and steady upward trend was observed in female suicide mortality, despite male suicide rates remaining consistently higher than female rates throughout the study period. The increase was particularly pronounced among younger age groups of females.
We also found that socioeconomic and contextual factors play a significant role in suicide rates. For males, suicide rates showed a statistically significant positive association with the proportion of the population living in rural areas, with a 10 percentage-point increase in rural population associated with an estimated 5.4% rise in male suicide mortality. For females, the annual unemployment rate was a significant factor, with a one percentage-point increase linked to a 2.4% rise in female suicide mortality. While regional unemployment rates were associated with suicide mortality, these rates primarily reflect broader economic conditions rather than individual employment status. Thus, the observed associations should be interpreted as indicative of the general socioeconomic context influencing suicide risk, rather than as direct causal effects of unemployment at the individual level.
Geographic disparities were also evident, with males exhibiting greater heterogeneity in rates across provinces, while females displayed a smoother spatial distribution. Despite these differences, two regions, Galicia in the northwest and southern Andalusia, were consistently identified as having suicide rates markedly above the national mean for both sexes. The study also concluded that age is a major determinant of suicide risk, with the age group component accounting for the largest proportion of total variability in all models, explaining 99% of the variability for males and 96% for females.
List of abbreviations
CrI: credible interval, DIC: Deviance Information Criterion, iCAR: intrinsic conditional autoregressive, ICD: International Classification of Diseases, INE: National Statistics Institute (Spain), INLA: integrated nested Laplace approximations, RSR: restricted spatial regression, RW1: first-order random walk, RW2: second-order random walk, WAIC: Watanabe-Akaike Information Criterion.
Declarations
Ethics approval and consent to participate
Not applicable.
Consent for publication
Not applicable.
Availability of data and materials
All analyses in this paper are fully reproducible. The data and R code for fitting the models with R-INLA, as well as for generating all tables and figures, are publicly available at: https://github.com/spatialstatisticsupna/Suicides_Spain.
Competing interests
The authors declare no competing interests.
Funding
@@ 6 Conclusion
This work is supported by projects PID2024-155382OB-I00 (funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE), PID2020-113125RB-I00/MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation), and UNEDPAM/PI/PR24/01A (Centro Asociado UNED - Pamplona).
Authors’ contributions
All authors contributed equally to this work.
Acknowledgements
Not applicable.